OpenStudy (anonymous):

Write an equation of the line containing the given point and parallel to the given line. (5,-6); 7x-9y=4

OpenStudy (anonymous):

First get &x-9x=4 to y=mx+b form

OpenStudy (anonymous):

7x-9y=4 into y=mx+b